OpenStudy (anonymous):

I need help with two problems Im just not getting them at all PLEASE!!! -8 – 52 ÷ 13 x 32 + 5 and 3 + 4[(8 ÷ 2) + 3]

OpenStudy (anonymous):

thats not 32 it is -8 -52\13*3^2+5

OpenStudy (phi):

If you want some more insight, here is a video by Khan http://www.khanacademy.org/video/order-of-operations?topic=developmental-math-1-1 (and the following 2)

OpenStudy (anonymous):

this is about order of operations. 1) start with anything inside parentheses 2) moving from left to right do exponents/multiplication/division FIRST, THEN move to addition/subtraction.

OpenStudy (anonymous):

-8-52/13x3^2 + 5 = -60/13x9 + 5 = -60/117+5 = -60/122 =-0.491 round off to -0.50 As for your second question: 3 + 4[(8 ÷ 2) + 3] 3 +4[(4) +3] 3 +4[7] 3 +28 31

OpenStudy (anonymous):

-8 -52/13*3^2+5 (so first, we'll do the squaring) -8 - 52 / 13 * 9 +5 (next, let's do the division) -8 - 4*9+5 (now, multiplication) -8 - 36 + 5 (no subtract and add, left to right) -44 + 5 ((finish it off) -39

OpenStudy (radar):

Memorize the order of operations. I used to know a mneomic that helped but forgot it lol

OpenStudy (anonymous):

3 + 4[(8 ÷ 2) + 3] (first, parentheses, innermost!) 3 + 4 [(4+3)] 3 + 4(7) 3+28 = 31

OpenStudy (anonymous):

how do I Evaluate 4 – 7x for x = -6 how do I do this ?

OpenStudy (anonymous):

just subsitute -6 in for everytime you see an X. so: 4 - 7(-6)

OpenStudy (anonymous):

4-7x So, 4-7(-6) 4+42=46

OpenStudy (anonymous):

and then you can solve (remember to do the parentheses first. 4 - (-42) 4+ 42 = 46

OpenStudy (anonymous):

see I never went to high school I ended up going to home school for my high school year and now Im in college and trying to learn all this math :( :(

OpenStudy (anonymous):

thank you all you have been a very big help to me :)

OpenStudy (anonymous):

do you understand how we worked the problems out?

OpenStudy (anonymous):

yes now I do a little bit better my only other one I cant figure out is (6÷ 3) + 4(6÷2)2

OpenStudy (anonymous):

that last two is ^2

OpenStudy (anonymous):

so would I divid first

OpenStudy (anonymous):

\[6\div3+4(6\div2)^{2}\] like that??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, so the first thing to look for are parentheses. this means you have to do the 6 / 2 FIRST! what is 6/2?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

good! so now we have this: 6÷3+4(3)^2 next, do the exponenet. 3^2 = ??

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Answer is: 2+4(3)^2 or 2 +4(9) = 2 + 36 = 38

OpenStudy (anonymous):

3^2 = 3*3 = ??

OpenStudy (anonymous):

9 right?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Yes 3^2 is square of 3.

OpenStudy (anonymous):

great! so now we have this: 6÷3+4(9) next, you work from LEFT to RIGHT.... do the division and mutliplication first... so what is 6/3 = ?? and what is 4*9 = ???

OpenStudy (anonymous):

6\3=2 4*9=36 right

OpenStudy (anonymous):

awesome. now you can put it all together. 2 + 36 = ?? and that's your answer!

OpenStudy (anonymous):

38

OpenStudy (anonymous):

great! did that make sense?

OpenStudy (anonymous):

yes it did thank you so much !!!!!!!!!

OpenStudy (anonymous):

no problem.